If a 4.0x10¹⁸ atoms decay with a life of 2.3 years, how many are remaining after 3.7 years?

1
Expert's answer
2022-06-03T15:25:31-0400
"N=N_0*2^{-t/t_{1/2}}\\
=4.0*10^{18}*2^{-3.7/2.3}=1.3*10^{18}"